# Unisolver: PDE-Conditional Transformers Are Universal PDE Solvers

### v3

@@ 3.2 Universal Components Embedding
Figure 2: Overview of Unisolver. We universally embed all PDE components into deep conditions and employ a conditional Transformer to aggregate deep conditions in the decoupled subspace.
As described in Section 3.1, PDE solutions are obtained by intricate interactions between initial conditions and complete equation components which can be grouped into two categories. In previous works (Brandstetter et al., 2022, Takamoto et al., 2023), these equation components are coarsely and incompletely included as conditions to modulate the input observations. In this paper, we will elaborate on how Unisolver finely and completely embeds all considered PDE components (Table 1) into deep PDE conditions based on our insights from the mathematical analysis.
@@
Since the mathematical symbols convey rich mathematical information, we utilize a Large Language Model (LLM) for symbolic embedding. Specifically, we adopt the recently released LLaMA-3 8B model  [^1] https://ai.meta.com/blog/meta-llama-3/ to embed the equation formulation. We attempt to leverage its understanding of prior mathematical knowledge, which was learned from pre-training on 15 TB of language tokens, as well as its flexible encoding of unstructured PDE information. Technically, the input to the LLM is the LaTeX code of the equation. For example, the Eq. (1a) is prompted as
@@
Then we take the output of the last Transformer block of the LLM and average representations along the sequence dimension, resulting in a 4096-dimensional embedding for each PDE. Notably, in the LLM embedding stage, we utilize mathematical symbols of the remaining equation components (e.g. coefficients and force terms) rather than their actual values in the prompt. For instance, we adopt symbol “a” in the above prompt rather than its concrete value to make the LLM focus on the key physics meaning of PDEs. The embedding of concrete values for the other components will be detailed in the next paragraph. After the LLM embedding stage, the hidden representations of PDE symbols are encoded by an MLP to align channel dimensions and obtain deep conditions.
@@
As we illustrated in Table 1, other components can be categorized as domain-wise and point-wise according to their effect on the final solution. Correspondingly, we adopt different embedding methods for these two types. For domain-wise components, including coefficients represented as real-valued vectors and boundary types similar to class labels, we embed them using two linear layers with an in-between SiLU activation function (Elfwing et al., 2018). Moreover, point-wise components like external force, binary geometry mask, and boundary value functions are essentially physical fields observed on mesh $\mathcal{M}$. We apply the same patchify embedding method used for input observations, transforming them into deep representation sequences.
@@
As shown in Figure 2, after universal components embedding, deep conditions within the same category are aggregated together to consolidate their impact. This strategy prevents excessive separation of deep PDE conditions that could weaken the model’s expressive capabilities, and thus will enhance representation learning for diverse PDE solving via joint training.
@@ 3.3 PDE-Conditional Transformer
We propose a conditional Transformer to adaptively fuse deep PDE conditions, embedded from the complete equation components, into hidden representations of inputs within decoupled subspaces.
@@
We evenly split the hidden representations of the inputs along the channel dimension, with one half influenced by domain-wise deep conditions and the other half by point-wise deep conditions. Especially in multi-head attention (Vaswani et al., 2017), our proposed subspace decoupling is equivalent to assigning some heads to learn the impact of domain-wise conditions while others focusing on point-wise conditions. This leads to improved representation learning for both categories, and minimized interference between deep PDE conditions from two categories.
@@
We utilize MLPs to individually project domain-wise conditions and point-wise conditions into the corresponding subspace. After projection, domain-wise conditions are repeated along the sequence dimension to match the length of token sequence and ensure consistent physical guiding throughout the entire sequence. The transformed conditions convey both domain-wise and point-wise information, which are then integrated adaptively by aggregation functions.
As shown in Figure 2, we aggregate conditions either before or after the attention and feedforward modules within Transformer. Inspired by recent conditional Transformers like DiT (Peebles & Xie, 2023) and other conditional normalization approaches (Park et al., 2019, Perez et al., 2018), we take the aggregation paradigm to finely capture the intricate correlations between hidden inputs of initial observations and deep equation conditions. Specifically, we scale and shift the hidden representations of inputs based on the equation conditions. After passing through the Transformer modules, we use the equation conditions to softly select whether this information should be retained.
@@
Summarizing the above designs, we propose the Unisolver (Figure 2). Given input $\mathbf{X}$, it is projected to embeddings $\mathbf{X}^{0}$ using a patchify layer (Dosovitskiy et al., 2020). The complete PDE equation components $\mathcal{C}_{\text{PDE}}$ are embedded into deep conditions $\mathbf{C}_{\text{domain}}$ and $\mathbf{C}_{\text{point}}$ following Section 3.2. Suppose there are $N$ layers, the $n$-th layer of Unisolver can be formalized as:
$\begin{split}\mathbf{I}_{{\ast}}&=\operatorname{Concat}\big{(}{\operatorname{MLP}_{\ast}}(\mathbf{C}_{\text{domain}}).\texttt{repeat},{\operatorname{MLP}_{\ast}}(\mathbf{C}_{\text{point}})\big{)},\quad\ast\in\{\text{scale, shift, select}\}\\ {\widehat{\mathbf{X}}}^{n-1}&=\mathbf{I}_{\text{select}}\odot\operatorname{SelfAttention}\big{(}\mathbf{I}_{\text{scale}}\odot\operatorname{{LayerNorm}}({\mathbf{X}}^{n-1})+\mathbf{I}_{\text{shift}}\big{)}+{\mathbf{X}}^{n-1},\\ {\widehat{\mathbf{I}}}_{\ast}&=\operatorname{Concat}\big{(}{{\widehat{\operatorname{MLP}}}_{\ast}}(\mathbf{C}_{\text{domain}}).\texttt{repeat},{{\widehat{\operatorname{MLP}}}_{\ast}}(\mathbf{C}_{\text{point}})\big{)},\!\!\quad\ast\in\{\text{scale, shift, select}\}\\ {\mathbf{X}}^{n}&={\widehat{\mathbf{I}}}_{\text{select}}\odot\operatorname{FeedForward}\big{(}{\widehat{\mathbf{I}}}_{\text{scale}}\odot\operatorname{LayerNorm}({\widehat{\mathbf{X}}}^{n-1})+{\widehat{\mathbf{I}}}_{\text{shift}}\big{)}+{\widehat{\mathbf{X}}}^{n-1},\end{split}$ (3)
where $n\in\{1,\ldots,N\}$, and ${\mathbf{X}}^{n}$ is the output of the $n$-th layer. Since the PDE components have a crucial impact on the range of the output, we scale and shift $\mathbf{X}^{N}$ based on the deep equation conditions, and then linearly project $\mathbf{X}^{N}$ to obtain the final output as predictions of $\mathbf{Y}$.
@@ 4 Experiments
We conduct extensive experiments to evaluate Unisolver on three challenging large-scale benchmarks, covering a wide range of PDE components and diverse generalization scenarios.
@@
As summarized in Table 2, three experimental large-scale benchmarks cover varied dimensions, resolutions and PDE components. The HeterNS is an extension of the NS dataset from FNO (2021a), incorporating multiple viscosity coefficients and external forces to enhance diversity. The 1D time-dependent PDEs, introduced by PDEformer (2024), is a large-scale dataset containing three million structured 1D PDE samples and evaluate the zero-shot generalization performance on PDEBench (2022). The 2D mixed PDEs, collected by DPOT (2024), include 12 diverse datasets from four well-established benchmarks. More details can be found in Appendix E.
Table 2: Summary of benchmarks. #GPU hours are calculated by averaging the training time of all models on one A100 GPU. Detailed compute resources can be found in Appendix G.6. $\checkmark$ indicates the PDE component will change among different samples, while $\times$ refers to unchanged ones.
Benchmarks #Dim #Resolution #Samples #GPU hours Symbols Coefficient Force Geometry Boundary
HeterNS 2D+Time (64,64,10) 15k $\sim$60h $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$
1D time-dependent PDEs 1D+Time (256,100) 3M $\sim$3000h $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
2D mixed PDEs 2D+Time (128,128,10) 74.1k $\sim$800h $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
We compare Unisolver with six advanced baselines on the HeterNS to demonstrate its generalizability under varied PDE components: the well-established FNO (2021a), PINO (2021b) and ViT (2020) and current state-of-the-art methods Factformer (2023b), ICON (2023) and MPP (2023). We augment these baselines by providing sufficient physics information to ensure a fair comparison, either by concatenating the inputs with varied PDE components, providing prompting trajectories (ICON) or applying physics-informed loss (PINO). Furthermore, we compare Unisolver with two generalizable solvers—PDEformer (2024) and DPOT (2024) on zero-shot generalization performance in a head-to-head manner. We refrain from including additional baselines on these two benchmarks due to the substantial computational cost of using million-scale samples.
@@
All methods in the HeterNS benchmark are trained for 300 epochs using relative L2 loss and the ADAM optimizer (Kingma & Ba, 2015) with an initial learning rate of $0.0005$ and a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2016). The batch size is set to 60. For the 1D time-dependent PDEs and 2D mixed PDEs, we follow the training strategies from PDEformer (2024) and DPOT (2024) to ensure a fair comparison. Relative L2 is used as the evaluation metric. See Appendix F for full implementation details and hyper-parameter configurations.
4.1 Heterogeneous 2D Navier-Stokes Equation (HeterNS)
Figure 3: Visualization of various evaluation scenarios on the HeterNS benchmark.
@@ 4.1 Heterogeneous 2D Navier-Stokes Equation (HeterNS)
We introduce HeterNS, an extension of the widely used 2D NS dataset (Li et al., 2021a), to assess how models handle diverse PDE components, particularly viscosity coefficients and force terms. It comprises five viscosity coefficients $\nu$ and three force terms differentiated by frequency $\omega$, resulting in 15 combinations of PDE components and 15,000 training samples. As depicted in Figure 3, we evaluate the model performance on in-distribution test with only unseen initial conditions and zero-shot generalization involving both unseen initial conditions and variations in viscosity coefficients or force terms.

### v4

@@ 3.2 Universal Components Embedding
As described in Section 3.1, PDE solutions are obtained by intricate interactions between initial conditions and complete equation components which can be grouped into two categories. In previous works [3, 38], these equation components are coarsely and incompletely included as conditions to modulate the input observations. In this paper, we will elaborate on how Unisolver finely and completely embeds all considered PDE components (Table 1) into deep PDE conditions based on our insights from the mathematical analysis.
@@
Since the mathematical symbols convey rich mathematical information, we utilize a Large Language Model (LLM) for symbolic embedding. Specifically, we adopt the LLaMA-3 8B model  [^1] https://ai.meta.com/blog/meta-llama-3/ to embed the equation formulation. We attempt to leverage its understanding of prior mathematical information, which was learned from pre-training on 15 TB of language tokens, as well as its flexible encoding of unstructured PDE information. Technically, the input to the LLM is the LaTeX code of the equation. For example, the Eq. (1a) is prompted as
@@
Then we take the output of the last Transformer block of the LLM and average the representations along the sequence dimension, resulting in a 4096-dimensional embedding for each PDE. In the LLM embedding stage, we use mathematical symbols of the remaining equation components (e.g., coefficients and force terms) instead of their actual values in the prompt. For instance, we use the symbol “a” in the prompt rather than its concrete value to make the LLM focus on the key physics meaning of the PDEs. The embedding of concrete values for the other components is detailed in the next paragraph. After the LLM embedding stage, the hidden representations of PDE symbols are encoded by an MLP to align channel dimensions and obtain deep conditions.
@@
As shown in Table 1, other components can be categorized as domain-wise and point-wise based on their effect on the final solution. Correspondingly, we adopt different embedding methods for these two types. For domain-wise components, including coefficients represented as real-valued vectors and boundary types akin to class labels, we embed them using two linear layers with an intermediate SiLU activation function [7]. For point-wise components like external force, binary geometry mask, and boundary value functions are essentially physical fields observed on mesh $\mathcal{M}$. We apply the same patchify embedding method used for input observations, transforming them into deep representation sequences.
@@
As shown in Figure 3, after universal components embedding, deep conditions within the same category are added together to consolidate their impact. This strategy prevents excessive separation of deep PDE conditions that could weaken the model’s expressive capabilities, and thus will enhance representation learning for diverse PDE solving via joint training.
@@ 3.3 PDE-Conditional Transformer
We propose a conditional Transformer to adaptively fuse deep PDE conditions which are embedded from the complete equation components, into hidden representations of inputs within decoupled subspaces.
@@
We evenly split the hidden representations of the inputs along the channel dimension, with one half influenced by domain-wise deep conditions and the other half by point-wise deep conditions. In multi-head attention [40], our proposed subspace decoupling is equivalent to assigning some heads to learn the impact of domain-wise conditions while others focusing on point-wise conditions. This leads to improved representation learning for both categories, and minimized interference between deep PDE conditions from two categories.
@@
We use MLPs to individually project domain-wise and point-wise conditions into corresponding subspaces. After projection, domain-wise conditions are repeated along the sequence dimension to match the token sequence length and ensure consistent physical guiding throughout the sequence. The transformed conditions convey both domain-wise and point-wise information and are integrated adaptively by aggregation functions.
As shown in Figure 3, we aggregate conditions either before or after the attention and feedforward modules within Transformer. Inspired by recent conditional Transformers like DiT [31] and other conditional normalization approaches [30, 32], we take the aggregation paradigm to finely capture the intricate correlations between hidden inputs of initial observations and deep equation conditions. Specifically, we scale and shift the hidden representations of inputs based on the equation conditions. After passing through the Transformer modules, we use the equation conditions to softly select whether this information should be retained.
@@
Summarizing the above designs, we propose the Unisolver (Figure 3). Given input $\mathbf{X}$, it is projected to embeddings $\mathbf{X}^{0}$ using a patchify layer [6]. The complete PDE equation components $\mathcal{C}_{\text{PDE}}$ are embedded into deep conditions $\mathbf{C}_{\text{domain}}$ and $\mathbf{C}_{\text{point}}$ following Section 3.2. Suppose there are $N$ layers, the $n$-th layer of Unisolver can be formalized as:
$\begin{split}\mathbf{I}_{{\ast}}&=\operatorname{Concat}\big{(}\operatorname{MLP}_{\ast}(\mathbf{C}_{\text{domain}}).\texttt{repeat},\operatorname{MLP}_{\ast}(\mathbf{C}_{\text{point}})\big{)},\\ {\widehat{\mathbf{X}}}^{n-1}&=\mathbf{I}_{\text{select}}\odot\operatorname{SA}\big{(}\mathbf{I}_{\text{scale}}\odot\operatorname{{LN}}({\mathbf{X}}^{n-1})+\mathbf{I}_{\text{shift}}\big{)}+{\mathbf{X}}^{n-1},\\ {\widehat{\mathbf{I}}}_{\ast}&=\operatorname{Concat}\big{(}{\widehat{\operatorname{MLP}}}_{\ast}(\mathbf{C}_{\text{domain}}).\texttt{repeat},{\widehat{\operatorname{MLP}}}_{\ast}(\mathbf{C}_{\text{point}})\big{)},\\ {\mathbf{X}}^{n}&={\widehat{\mathbf{I}}}_{\text{select}}\odot\operatorname{FF}\big{(}{\widehat{\mathbf{I}}}_{\text{scale}}\odot\operatorname{LN}({\widehat{\mathbf{X}}}^{n-1})+{\widehat{\mathbf{I}}}_{\text{shift}}\big{)}+{\widehat{\mathbf{X}}}^{n-1},\end{split}$ (2)
[…]
@@ 4 Experiments
We conduct extensive experiments to evaluate Unisolver on three challenging large-scale benchmarks, covering a wide range of PDE components and generalization scenarios.
@@
As summarized in Table 2, three experimental large-scale benchmarks cover varied dimensions, resolutions and PDE components. The HeterNS is an extension of the NS dataset from FNO [2021a], incorporating multiple viscosity coefficients and external forces to enhance diversity. The 1D time-dependent PDEs, introduced by PDEformer [2024], is a large-scale dataset containing three million structured 1D PDE samples and evaluate the zero-shot generalization performance on PDEBench [2022]. The 2D mixed PDEs, collected by DPOT [2024], include 12 diverse datasets from four well-established benchmarks. More details can be found in Appendix G.
Table 2: Summary of benchmarks. #GPU hours are calculated by averaging the training time of all models on one A100 GPU. Detailed compute resources can be found in Appendix I.7. Besides, $\checkmark$ indicates the PDE component will change among different samples, while $\times$ refers to unchanged ones.
Benchmarks HeterNS 1D time-dependent PDEs 2D mixed PDEs
#Dim 2D+Time 1D+Time 2D+Time
#Resolution (64,64,10) (256,100) (128,128,10)
#Samples 15k 3M 74.1k
#GPU hours $\sim$60h $\sim$3000h $\sim$800h
Symbols $\times$ $\checkmark$ $\checkmark$
Coefficient $\checkmark$ $\checkmark$ $\checkmark$
Force $\checkmark$ $\checkmark$ $\checkmark$
Geometry $\times$ $\times$ $\checkmark$
Boundary $\times$ $\checkmark$ $\checkmark$
@@
We compare Unisolver with six advanced baselines on the HeterNS to demonstrate its generalizability under varied PDE components: the well-established FNO [2021a], PINO [2021b] and ViT [2020] and current state-of-the-art methods FactFormer [2023b], ICON [2023] and MPP [2023]. We augment these baselines by providing sufficient physics information to ensure a fair comparison, either by concatenating inputs with varied PDE components, providing prompting trajectories (ICON), or applying physics-informed loss (PINO).
Furthermore, we compare Unisolver with two generalizable solvers—PDEformer [2024] and DPOT [2024] on zero-shot generalization performance. We refrain from including additional baselines on these two benchmarks due to the substantial computational cost of using million-scale samples.
@@
All methods in the HeterNS benchmark are trained for 300 epochs using relative L2 loss and the ADAM optimizer [17] with an initial learning rate of $0.0005$ and a cosine annealing learning rate scheduler [26]. The batch size is set to 60. For the 1D time-dependent PDEs and 2D mixed PDEs, we follow the training strategies from the original papers of PDEformer [2024] and DPOT [2024] to ensure a fair comparison. Relative L2 is used as the evaluation metric. See Appendix H for full implementation details and hyper-parameter configurations of each model.
4.1 Heterogeneous 2D Navier Stokes Equation: HeterNS
@@
We introduce HeterNS, an extension of the widely used 2D NS dataset [19], to assess how models handle diverse PDE components, particularly viscosity coefficients and force terms. It comprises five viscosity coefficients $\nu$ and three force terms differentiated by frequency $\omega$, resulting in 15 combinations of PDE components and 15,000 training samples. As depicted in Figure 4, we evaluate the model performance on in-distribution test with only unseen initial conditions and zero-shot generalization involving both unseen initial conditions and variations in viscosity coefficients or force terms.
Figure 4: Visualization of evaluation scenarios on the HeterNS.
Table 3: Performance comparison on HeterNS. Note that the reported In-dist and Zero-shot results are averaged over multiple sub-configurations with different viscosity coefficients and force terms. Full results can be found in Appendix D.
Method #Params Viscosity Force
In-dist Zero-shot In-dist Zero-shot
FNO 4.7M 0.0210 0.0574 0.0975 0.4316
PINO 4.7M 0.0364 0.0679 0.1544 0.7155
ViT 4.8M 0.0156 0.0514 0.0593 0.3158
FactFormer 5.1M 0.0201 0.1162 0.0654 0.3272
ICON 4.5M 0.0216 0.0611 0.0788 0.7621
MPP 4.9M 0.0347 0.0840 0.0997 0.3435
Unisolver 4.1M 0.0098 0.0374 0.0428 0.1053
Promotion / 37.2% 27.2% 27.7% 66.7%
1D Time-dep endent PDEs Params Tasks In-distribution Test
Zero-shot Burgers
Zero-shot Advection
$\nu=0.1$ $\nu=0.01$ $\nu=0.001$ $\beta=0.1$
PDEformer 22M 0.0225 0.00744 0.0144 0.0393 0.0178
Unisolver 19M 0.0108 0.00513 0.00995 0.0299 0.0138
Promotion / 52.0% 31.0% 30.9% 23.9% 22.5%
Table 5: Performance comparison (relative L2 (×$\times 10^{-2}$)) across 12 in-distribution test sets. We use a “source-PDE” format to denote different tasks (e.g. FNO-NS). The second row lists the primary PDE components for each dataset. See Appendix G.3 for details.
2D Mixed PDEs Params Tasks
FNO-NS-$\nu$
PDEBench-CNS-(M,$\zeta$)
PDEBench
PDEArena
CFDBench-NS
Average Error
1e-5
1e-4
1e-3
(1, 0.1)
(1, 0.01)
(0.1, 0.1)
(0.1, 0.01)
DR
SWE
NS
NS-Force
Geometry
DPOT
30M
5.53
4.42
1.31
1.53
3.37
1.19
1.87
3.79
0.66
9.91
31.6
0.70
5.50
Unisolver
33M
4.17
3.36
0.61
1.23
2.89
1.01
1.59
4.39
0.45
6.87
27.4
0.54
4.54
Promotion (%)
24.6
24.0
53.4
19.6
14.2
15.1
15.0
31.8
30.7
13.3
22.9
17.5